what is the lewis dot structure of XeF2O? Draw all lone pairs.
Answer #1
Xe would be your central atom, O is attached to it by a double bond and your two Fs are attached be a single bond. There would be 3 Lone Pair (LP) electrons around F, 2 around O, and 2 around Xe. Why ? There is no valance to this molecule so you want to satisfy the formal charge to be 0. Formal charge = (#electrons in ground state) - [ (free electrons) - 1/2 (bonding electrons)]. Formal charge for F, would want 0 so 0 = 7 - [? + 1/2 (2) ] = 7 - (? +1) = 7 - ? - 1 ? = 6, therefore 3 lone pairs around fluorines. Formal charge for O, would want 0 so 0 = 6 - [? + 1/2 (4) ]... ? = 4, therefore 2 LP around oxygen Formal charge for Xe, again we want it to be 0 so, 0 = 8 - [? + 1/2 (8)]..... ? = 4, therefore 2 LP around Xe.